# FakET: Simulating Cryo-Electron Tomograms with Neural Style Transfer

### v2

@@ Abstract
Particle localization and -classification constitute two of the most fundamental problems in computational microscopy. In recent years, deep learning based approaches have been introduced for these tasks with great success. A key shortcoming of these supervised learning methods is their need for large training data sets, typically generated from particle models in conjunction with complex numerical forward models simulating the physics of transmission electron microscopes. Computer implementations of such forward models are computationally extremely demanding and limit the scope of their applicability. In this paper we propose a method for simulating the forward operator of an electron microscope based on additive noise and Neural Style Transfer techniques. We evaluate the method on localization and classification tasks using one of the established state-of-the-art architectures showing performance on par with the benchmark. In contrast to previous approaches, our method accelerates the data generation process by a factor of 750 while using 33 times less memory and scales well to typical transmission electron microscope detector sizes. It utilizes GPU acceleration and parallel processing. It can be used to adapt a synthetic training data set according to reference data from any transmission electron microscope. The source code is available at https://gitlab.com/deepet/faket.
@@ 1 Introduction
Model Training data Data cost Localization F1 Classification F1
@@
Table 1: Comparison of DeepFinder’s performance on localization and classification tasks as a function of training data. The performance is evaluated on the same test tomogram. The results are shown in context with the performance of standard template matching algorithms (denoted TM and TM-F) reported in Gubins et al. (2021) for the same testing data. Performance is measured using the F1 score for localization and the F1 macro score for classification. Each of the DF results is an average of the test performances measured at the best epoch (based on validation) over 6 different random seeds. On the challenging classification task, DF trained using faket data simulated by our proposed method reaches 92 % of the performance of DF trained using benchmark data. It even reaches 97 % when fine-tuned using a portion of benchmark data. All this for a fraction of computational cost and without the need for a configuration protocol of the original simulation parameters used for crating benchmark.
Recent developments in cryo-electron tomography (cryoET) allow to obtain high resolution representations of macromolecular complexes in their native cellular environment showing molecular interactions that are hardly accessible with other methods Turk and Baumeister (2020). In cryoET, the imaged sample is in most cases a 100-200 nm thick slice of a frozen cell. From this slice, projection images are taken in a transmission electron microscope (TEM) from different rotation (tilt) angles. An artifact free reconstruction would require measurements using tilt angles, that would complete the half circle. However, this is not feasible, due to limitations of the specimen holder, and only a range of 140 can be recorded. The missing tilt images later on result in a so-called missing-wedge in the 3D reconstruction (cryo-electron tomogram). In addition, the electron beam severely damages the sample during imaging, so only a low electron dose can be used to image a biological specimen. The low dose in combination with the presence of ice in the sample results in the acquired data being very noisy. Consequently, the identification of molecules within these reconstructions is a daunting task. Particle identification is however necessary as the particles need to be classified and averaged to determine high resolution structures. While cryoET has led to a large number of breakthroughs, providing hitherto unseen detail in the molecular architecture of cells (Zimmerli et al., 2021; O’Reilly et al., 2020; Mahamid et al., 2016), the aforementioned challenges still hinder the widespread use of cryoET in the larger cell biology and structural biology community. In this context, the development of new reliable software tools is of paramount importance, which is however obstructed by the lack of sufficient accessible and annotated data to develop the software tools on.
@@ 1.1 SHREC simulator
Unfortunately, simulating the tomographic data using SHREC is computationally very expensive. For a set of 10 small tilt-series (61 tilts of size $1024\times 1024$) it took approx. 150 CPU hours of computation on a node with $2\,\times$ Intel Xeon E5-2630 v4 CPUs. The implementation is able to utilize only 3 CPU cores per job in parallel and each job needs 114 GB of RAM. The memory consumption is also a reason why utilizing GPUs and simulating tilt-series of common sizes is not yet feasible (personal communication with the authors Gubins, I., and Chaillet, M.). This limits the scope of its applicability, mainly in data-hungry applications such as deep learning. Moreover, at the time of writing of this article, the source code of the simulator is not publicly available.
@@ 1.2 Our contribution
In this paper, we propose faket, a novel method for simulating the forward operator of a cryo transmission electron microscope to generate synthetic tilt-series. It was created, among other reasons, to generate training data for deep neural networks to solve tasks such as particle localization and (much more challenging) particle classification. Our method combines additive noise and neural style transfer (NST) techniques based on Gatys et al. (2016) and is evaluated on downstream tasks using DeepFinder (DF)  Moebel et al. (2021) – a neural network specifically tailored to the task of particle localization and classification, see also Section 4.1.
@@
faket is a surrogate model that mimics the behavior of the physics-based TEM simulator SHREC, while being drastically less computationally expensive, both in terms of time and memory (see Table 1). It also does not require any calibration protocol as other simulators, see Section 1.4. Moreover, it does not need to be retrained to be used on new data, nor does it require labeled reference data. In contrast to SHREC, our method accelerates the data generation process by a factor of 750 while using 33 times less memory. It therefore has the potential to save experts countless hours  [^1] Depending on the imaged sample, an expert may spend several hours of manual work per tilt-series per particle labeling the data. At the same time, certain smaller particles cannot be visually found at all, thus fully labeled tomograms do not really exist. of manual work in labeling their data sets.
Moreover, our method is capable of simulating large tilt-series, which are common in experimental environments (about $12\times$ larger than SHREC projections). For example, we generated a $61\times 3500\times 3500$ tilt-series on a single NVIDIA A100 40GB SXM4 GPU in less than 10 minutes. This advance makes it possible to train particle localization and classification networks from scratch, or to pre-train networks that are later fine-tuned using manually labelled experimental data, see Appendix C. On top of that, our method is open source and our experiments are reproducible.
@@
Additionally, we include comparisons of the benchmark and the proposed faket method with a simple baseline method based on additive Gaussian noise, which will be useful for practitioners who need to decide which method fits into their computational budget. The contribution of our method is further supported by an ablation study presented in Appendix D, where we also offer insight into the potential limits of DF on the studied data by conducting experiments using completely noiseless simulated tomograms.
@@ 1.5 Structure of this article
To build our proposed method, presented in Section 3 (further referred to as faket), we used a data-driven approach (i.e. no calibration protocol is needed). More information about the data we used is in Section 2. With our method, we managed to successfully approximate the input-output behavior, on par with the SHREC simulator (further referred to as benchmark), but for a fraction of its computational cost. To experimentally prove the need for our method, we also compared it to a simple addition of Gaussian noise (further referred to as baseline) that was naturally very fast to compute, but did not lead to comparable results on the downstream tasks.
@@ 2 Data
As stated earlier, we used the full-resolution grandmodels from the SHREC data set to create our own noiseless projections using Radon transform, as implemented in the scikit-image library for image processing in Python Van der Walt et al. (2014). We computed our own noiseless projections for three primary reasons. Firstly, SHREC only supplied full-resolution noiseless projections embedded within the simulated ice layer, a feature we also aimed to incorporate in our surrogate. Secondly, we intended to provide access to all implementation steps within the code-base. Lastly, it was necessary to produce noiseless reconstructions that could later be used to estimate the performance boundaries of DF on this data set. We also had to omit the usage of the reconstructions provided by SHREC because the exact configuration of all the steps is not published and the version of the PyTom software package used to create this data is not public. Also, the public version is not yet properly documented or straightforward to use. Thus, the only feasible option to ensure comparable results across our experiments was to create our own reconstructions from all the newly created projections described in Section 3. Nevertheless, we put a lot of effort into matching our reconstructions with the ones originally published in SHREC. We created them using the RadonTea Python package Müller, Paul (2013) using a custom filter described in Appendix B. For additional flexibility, we also implemented support for tomographic reconstructions using the well-known IMOD package Mastronarde and Held (2017). To ensure comparability also with previously published results, we used the original SHREC model_9 reconstruction as a test tomogram for all presented methods as it was done in all the past challenges.
@@ 3.1 Additive noise (baseline)
Figure 5: Diagram of steps to simulate the benchmark, baseline, faket, and noiseless projections and reconstructions. Red arrows highlight steps to reproduce SHREC data from which we use the last tomogram for testing. All methods except SHREC were filtered using a reverse-engineered filter (see Appendix B) because the SHREC filtering step is under-documented. The style projections never feature the same contents as the simulated ones (see Appendix A).
@@ 3.2 Neural Style Transfer (faket)
NST was built to render the semantic content of natural images in different styles. At its core lies VGG net Simonyan and Zisserman (2014), a convolutional neural network optimized for object recognition and -localization. Within the NST framework, the VGG net is used for extracting the content and style representations of so-called “content” and “style” images, which are provided as inputs. NST then iteratively updates the output image to simultaneously match the content representation and the style representation of the provided inputs at multiple scales. NST is described in the aforementioned paper in great detail and we encourage the reader to consult it if any questions would arise.
@@ 6 Conclusions
In this paper, we proposed faket, a method for simulating the forward operator of an electron microscope based on additive noise and Neural Style Transfer. The proposed method can be used for generating synthetic cryo-electron tilt-series which can be used to train deep neural networks to solve tasks such as particle localization and (much more challenging) particle classification. The field of cryo-electron tomography currently suffers from the lack of sufficient amounts of annotated data, and the proposed method aims to solve this problem. faket is a surrogate model that mimics the behavior of the physics-based TEM simulator SHREC, while being drastically less computationally expensive, both in terms of time and memory. It accelerates the data generation process by a factor of 750 while using 33 times less memory, making the generation of thousands of tilt-series feasible. Moreover, it is capable of simulating large tilt-series, which are common in experimental environments. For example, we generated a $61\times 3500\times 3500$ tilt-series on a single NVIDIA A100 40GB SXM4 GPU in less than 10 minutes. It also does not require any calibration protocol, it does not need to be retrained to be used on new data, nor does it require labeled reference data. The method is open source, and the experiments are reproducible.
@@
The quality of our approximations was evaluated using the DeepFinder network, which emerged from the SHREC challenge as one of the most successful. The results showed that the performance of models trained using our approximations is on par with the benchmark method on localization task and reached 92 % of its performance on classification task while significantly outperforming standard template matching algorithms. When further fine-tuned using a portion of benchmark data, the classification performance was improved to 97 %. This advancement simplifies the generation of fully-labeled, high-quality synthetic tilt-series that resemble experimental TEM data requiring analysis. This simulated data can either be used to train particle localization and classification neural networks from scratch, or serve as fpre-training data for networks that will be fine-tuned with manually labeled experimental data later. Researchers investigating reconstruction algorithms can also benefit from our simulator since the availability of ground truth allows for effortless evaluation of their novel methods.
@@
In future work, we will focus on further validation of our method using experimental data from TEM and we hope this advancement will serve as the basis for more method development in cryoET. Additionally, we plan to improve the method by using state-of-the-art neural networks and we want to provide the community of practitioners with a fully functional and easy-to-use piece of software to generate their synthetic data sets based on chosen particles from Protein Data Bank. The goal is to enable more accurate and efficient data analysis while also making the process more accessible to researchers in the field.

### v3

@@ Abstract
In cryo-electron microscopy, accurate particle localization and classification are imperative. Recent deep learning solutions, though successful, require extensive training data sets. The protracted generation time of physics-based models, often employed to produce these data sets, limits their broad applicability. We introduce FakET, a method based on Neural Style Transfer, capable of simulating the forward operator of any cryo transmission electron microscope. It can be used to adapt a synthetic training data set according to reference data producing high-quality simulated micrographs or tilt-series. To assess the quality of our generated data, we used it to train a state-of-the-art localization and classification architecture and compared its performance with a counterpart trained on benchmark data. Remarkably, our technique matches the performance, boosts data generation speed $\bm{750\times}$, uses $\bm{33\times}$ less memory, and scales well to typical transmission electron microscope detector sizes. It leverages GPU acceleration and parallel processing. The source code is available at https://github.com/paloha/faket/.
@@
Author Contribution— Conceptualization, P.H. and D.H.; Methodology, Software, Validation, Formal Analysis, Investigation, P.H. and L.H.; Data Curation, and Visualization, P.H.; Resources, P.G.; Writing - Original Draft, P.H., P.G., and D.H.; Writing - Review & Editing, P.H. and D.H.; Supervision, P.G. and D.H.; Funding Acquisition, P.G. and D.H.
Declaration of Interests— The authors declare no competing interests.
@@ 1 Introduction
Recent developments in cryo-electron tomography (cryoET) allow to obtain high resolution representations of macromolecular complexes in their native cellular environment showing molecular interactions that are hardly accessible with other methods Turk and Baumeister (2020).
Model Training data Data cost Localiz. F1 Classification F1
@@
Table 1: Evaluation of simulated data quality by examining DeepFinder’s performance on localization and classification tasks as a function of training data. The performance is evaluated on the same test tomogram. The results are shown in context with the performance of standard template matching algorithms (denoted TM and TM-F) reported in Gubins et al. (2021) for the same testing data. Performance is measured using the F1 score for localization and the F1 macro score for classification. Each of the DF results is an average of the test performances measured at the best epoch (based on validation) over 6 different random seeds. On the challenging classification task, DF trained using faket data simulated by our proposed method reaches 92 % of the performance of DF trained using benchmark data. It even reaches 97 % when fine-tuned using a portion of benchmark data. All this for a fraction of computational cost and without the need for a configuration protocol of the original simulation parameters used for crating benchmark. The cost of data is given for the whole data set comprising 10 tilt-series of shape ($61\times 1024\times 1024$), i.e. generating one such tilt-series with faket takes $\approx 70$ seconds assuming a GPU is available. For comparison, a usual-sized tilt-series of shape ($61\times 3500\times 3500$), that is $\approx 12\times$ larger in comparison to benchmark, can be simulated under 10 minutes. For a comprehensive time and memory consumption profiling of faket, see Appendix F.
In cryoET, the imaged sample is in most cases a 100-200 nm thick slice of a frozen cell. From this slice, projection images are taken in a transmission electron microscope (TEM) from different rotation (tilt) angles. An artifact free reconstruction would require measurements using tilt angles, that would complete the half circle. However, this is not feasible, due to limitations of the specimen holder, and only a range of 140 can be recorded. The missing tilt images later on result in a so-called missing-wedge in the 3D reconstruction (cryo-electron tomogram). In addition, the electron beam severely damages the sample during imaging, so only a low electron dose can be used to image a biological specimen. The low dose in combination with the presence of ice in the sample results in the acquired data being very noisy. Consequently, the identification of molecules within these reconstructions is a daunting task. Particle identification is however necessary as the particles need to be classified and averaged to determine high resolution structures. While cryoET has led to a large number of breakthroughs, providing hitherto unseen detail in the molecular architecture of cells (Zimmerli et al., 2021; O’Reilly et al., 2020; Mahamid et al., 2016), the aforementioned challenges still hinder the widespread use of cryoET in the larger cell biology and structural biology community. In this context, the development of new reliable software tools is of paramount importance, which is however obstructed by the lack of sufficient accessible and annotated data to develop the software tools on.
@@ 1.1 SHREC simulator
Unfortunately, simulating the tomographic data using SHREC is computationally very expensive. For a set of 10 small tilt-series (61 tilts of size $1024\times 1024$) it took approx. 450 CPU hours of computation on a node with $2\,\times$ Intel Xeon E5-2630 v4 CPUs. The implementation is able to utilize only 3 CPU cores per job in parallel and each job needs 114 GB of RAM. The memory consumption is also a reason why utilizing GPUs and simulating tilt-series of common sizes is not yet feasible (personal communication with the authors Gubins, I., and Chaillet, M.). This limits the scope of its applicability, mainly in data-hungry applications such as deep learning. Moreover, at the time of writing of this article, the source code of the simulator is not publicly available prohibiting more detailed comparisons, e.g. in terms of FLOPs.
@@ 1.2 Our contribution
In this paper, we propose faket, a fast and scalable data-driven method for simulating the forward operator of any cryo transmission electron microscope with the aim to generate synthetic micrographs or tilt-series. It was created, among other reasons, to generate fully-labeled real-like data for training deep neural networks to solve tasks such as particle localization and (much more challenging) particle classification. Our method combines additive noise and neural style transfer (NST) technique based on Gatys et al. (2016) to capture, using only unlabeled reference data, the structure of the complex noise introduced by TEM. To carry out the NST, faket utilizes a pretrained model, eliminating the need for users to train it themselves. That means users can use faket to simulate data from any TEM and under any configuration, assuming they possess unlabeled TEM data that could serve as a style reference. faket delivers data of quality nearly identical [^1] In terms of their practical utility for subsequent tasks, rather than exact manifestation of the physical interactions between the sample and the transmission electron microscope. to the reference, at a cost only slightly higher than simply adding Gaussian noise to a noiseless synthetic sample [^2] Created using existing models of biological macromolecular structures, represented as Coulomb density volume.. This makes it a practical, effective, and efficient choice for simulation.
For the purpose of evaluating our method, in the experiments presented within this article, we employed faket to mimic the behavior of the physics-based TEM simulator SHREC [^3] Due to the availability of ground truth.. On top of evaluating the quality of our simulated data using standard image metrics (available in the supplement) that are of limited value in this context, we evaluated it directly on practically-relevant downstream tasks by training DeepFinder (DF)  Moebel et al. (2021) – a neural network specifically tailored to the task of particle localization and classification, see also Section 4.1.
@@
In contrast to SHREC, our method accelerates the data generation process by a factor of 750 while using 33 times less memory (see Table 1). It also does not require any calibration protocol as other simulators, see Section 1.4. The NST model does not need to be retrained to be used on new data, nor does it require labeled reference data. faket therefore has the potential to save experts countless hours [^4] Depending on the imaged sample, an expert may spend several hours of manual work per tilt-series per particle labeling the data. At the same time, certain smaller particles cannot be visually found at all, thus fully labeled tomograms do not really exist. of manual work in labeling their data sets. Moreover, faket is capable of simulating large tilt-series, which are common in experimental environments (about $12\times$ larger than SHREC projections). For example, we generated a $61\times 3500\times 3500$ tilt-series on a single NVIDIA A100 40GB SXM4 GPU in less than 10 minutes [^5] To achieve the reported speeds, a minimum of one GPU is recommended; CPU-only simulations are feasible and can be a valid option in specific situations, albeit slower. This advance makes it possible to train particle localization and classification networks from scratch, or to pre-train networks that are later fine-tuned using manually labelled experimental data, see Appendix C. On top of that, our method is open source and our experiments are reproducible.
@@
Additionally, we include comparisons of the benchmark and the proposed faket method with a simple baseline method based on additive Gaussian noise aiming at providing comparative results in full practical range. The costly SHREC method defines an upper-boundary of performance, while the cheap addition of Gaussian noise establishes a lower boundary. This configuration enables us to position our method along this spectrum, showcasing faket’s capability to offer results close to the upper-boundary while maintaining computational efficiency of a much simpler method. We hope our comparisons will be useful for practitioners who need to decide which method fits into their computational budget and for those seeking insightful understanding of the inherent trade-offs. The contribution of our method is further supported by an ablation study presented in Appendix D, where we also offer insight into the potential limits of DF on the studied data by conducting experiments using completely noiseless simulated tomograms.
@@ 1.5 Structure of this article
To build our proposed method, presented in Section 3 (referred to as faket), we used a data-driven approach (i.e. no calibration protocol is needed). More information about the data we used is in Section 2. With our method, we managed to successfully approximate the input-output behavior, on par with the SHREC simulator (referred to as benchmark), but for a fraction of its computational cost. To experimentally prove the need for our method, we also compared it to a simple addition of Gaussian noise (referred to as baseline) that was naturally very fast to compute, but did not lead to comparable results on the downstream tasks.
@@
1.6 Key findings
faket successfully utilizes NST to simulate cryo-electron micrographs or tilt-series of common sizes. • faket produces realistic TEM simulations of quality nearly identical to the unlabeled reference data. • Cost of faket simulation is only slightly higher than simply adding Gaussian noise while significantly lower than that of complex physics-based simulators. • Apart from unlabeled reference data, faket does not require further training or calibration. • faket is a practical and efficient tool for simulating fully-labeled training data for deep learning.
@@ 2 Data
As stated earlier, we used the full-resolution grandmodels from the SHREC data set to create our own noiseless projections using Radon transform, as implemented in the scikit-image library for image processing in Python Van der Walt et al. (2014). We computed our own noiseless projections for three primary reasons. Firstly, SHREC only supplied full-resolution noiseless projections embedded within the simulated ice layer, a feature we also aimed to incorporate in our surrogate. Secondly, we intended to provide access to all implementation steps within the code-base. Lastly, it was necessary to produce noiseless reconstructions that could later be used to estimate the performance boundaries of DF on this data set. We also had to omit the usage of the reconstructions provided by SHREC because the exact configuration of all the steps is not published and the version of the PyTom software package used to create this data is not public. Also, the public version is not yet properly documented or straightforward to use. Thus, the only feasible option to ensure comparable results across our experiments was to create our own reconstructions from all the newly created projections described in Section 3. Nevertheless, we put a lot of effort into matching our reconstructions with the ones originally published in SHREC. We created them using the RadonTea Python package Müller (2013) using a custom filter described in Appendix B. For additional flexibility, we also implemented support for tomographic reconstructions using the well-known IMOD package Mastronarde and Held (2017). To ensure comparability also with previously published results, we used the original SHREC model_9 reconstruction as a test tomogram for all presented methods as it was done in all the past challenges.
@@
3 Methods
@@ 3.1 Additive noise (baseline)
Figure 5: Diagram of steps to simulate the benchmark, baseline, faket, and noiseless projections and reconstructions. Red arrows highlight steps to reproduce SHREC data from which we use the last tomogram for testing. All methods except SHREC were filtered using a reverse-engineered filter (see Appendix B) because the SHREC filtering step is under-documented. The style projections never feature the same contents as the simulated ones (see Appendix A). Grandmodels, noiseless artificial samples containing randomly scattered particles, were created using existing models of biological macromolecular structures, represented as Coulomb density volumes.
@@ 3.2 Neural Style Transfer (faket)
NST was built to render the semantic content of natural images in different styles. At its core lies VGG net Simonyan and Zisserman (2014), a convolutional neural network optimized for object recognition and -localization. Within the NST framework, the VGG-19 model is used for extracting the content and style representations of so-called “content” and “style” images, which are provided as inputs. NST then iteratively updates the output image to simultaneously match the content representation and the style representation of the provided inputs at multiple scales. The NST technique is described in the aforementioned paper in great detail and we encourage the reader to consult it if any questions would arise. The VGG-19 model used in faket was pretrained on ImageNet data set of natural images, eliminating the need for users to train the model themselves. For those interested in further fine-tuning the NST model (benefits of which are still under investigation by the authors), guidelines are available on the PyTorch website. Fine-tuning would necessitate partially labeled data, unless an autoencoder approach is adopted, which would require no labels.
@@ 6 Discussion
In this paper, we proposed faket, a fast and scalable data-driven method for simulating the forward operator of any electron microscope based on additive noise and Neural Style Transfer. The proposed method can be used for generating synthetic cryo-electron micrographs or tilt-series that closely approximate the quality of reference TEM data, at a computational cost only marginally higher than that of simply adding Gaussian noise. The data can be used to train deep neural networks to solve tasks such as particle localization and (much more challenging) particle classification. The field of cryo-electron tomography currently suffers from the lack of sufficient amounts of annotated data, and the proposed method aims to solve this problem.
In this study, we evaluated faket as a surrogate model that mimics the behavior of the physics-based TEM simulator SHREC, while being drastically less computationally expensive, both in terms of time and memory. It accelerates the data generation process by a factor of 750 while using 33 times less memory, making the generation of thousands of tilt-series feasible. Moreover, it is capable of simulating large tilt-series, which are common in experimental environments. For example, we generated a $61\times 3500\times 3500$ tilt-series on a single NVIDIA A100 40GB SXM4 GPU in less than 10 minutes. It also does not require any calibration protocol, it does not need to be retrained to be used on new data, nor does it require labeled reference data. The method is open source, and the experiments are reproducible.
@@
The quality of our approximations was evaluated using the DeepFinder network, which emerged from the SHREC challenge as one of the most successful. The results showed that the performance of models trained using our approximations is on par with the benchmark method on localization task and reached 92 % of its performance on classification task while significantly outperforming standard template matching algorithms. When further fine-tuned using a portion of benchmark data, the classification performance was improved to 97 %.
This advancement simplifies the generation of fully-labeled, high-quality synthetic tilt-series that resemble experimental TEM data requiring analysis. This simulated data can either be used to train particle localization and classification neural networks from scratch, or serve as pre-training data for networks that will be fine-tuned with manually labeled experimental data later. Researchers investigating reconstruction algorithms can also benefit from our simulator since the availability of ground truth allows for effortless evaluation of their novel methods. However, when simulating data for use cases other than training neural networks for particle localization or classification, it is important to take into account that no explicit knowledge of the underlying physics phenomena of TEM is built into the simulator.
@@
In future work, we will focus on further validation of our method using experimental TEM data. Additionally, we plan to improve the method by replacing the VGG-19 network that was pretrained on natural images, with a state-of-the-art vision network pretrained on cryoEM data. Moreover, we aim to provide the community of practitioners with a fully functional and easy-to-use piece of software for generating their synthetic data sets. Either based on chosen particles from Protein Data Bank, or using already available whole-cell models. The goal is to enable more accurate and efficient data analysis while also making the process more accessible to researchers in the field. We hope this advancement will serve as the basis for development of new computational methods in cryoEM and cryoET.